# The Name of the Title Is Hope

### v2

@@ 5. Methodology
LightCL, as shown in Figure  3, efficiently delays forgetting through Maintaining Generalizability and Memorize Feature Patterns. The whole pipeline of them is shown in Figure  LABEL:fig:pipeline_. All these methods are based on the discovery of generalizability in Section 4. We first introduce Maintaining Generalizability in Section 5.1 and then illustrate Memorize Feature Patterns in Section 5.2.
Figure 3. Illustration of each component in LightCL. Maintaining Generalizability and Memorize Feature Patterns work together to overcome catastrophic forgetting on resource-constrained edge conditions.
5.1. Maintaining Generalizability
@@
According to the conclusion (Section 4) that lower and middle layers have strong generalizability, the training consumption on them can be a kind of redundancy. Hence, we easily freeze them to maintain generalizability without the training process, illustrated in the gray background section at the top of the Figure  3. Specifically, we must pretrain a network, freeze its lower and middle layers, and finally deploy it on the edge device for CL. Training the t-th task can be formulated as
@@
(3) $\begin{array}[]{cl}\theta_{i}^{t}\leftarrow\theta_{i}^{t}+\alpha\nabla_{\theta_{i}^{t}}\mathcal{L}_{c}(f(\theta^{t},x_{t}),y_{t})\quad\text{s.t.}\quad i>FZ,\end{array}$
@@
where FZ is the last layer index of frozen layers, $\mathcal{L}_{c}$ is the cross-entropy loss function, and function $f$ represents forward pass.
The frozen layers don’t involve backpropagation, so it does not need to store and compute many feature maps and gradients during training. We not only reduce memory footprint but also decrease training computation. As shown in Section 6.2.1, with the help of maintaining generalizability, LightCL  reduces at $6.16$$\times$ memory footprint and FLOPs are competitive beyond efficient CL  method SparCL. Meanwhile, in Section 6.3, it has greatly delayed forgetting in most CL  methods, which proves the universality of our method.
@@ 5.2.1. How to Memorize Efficiently
Forgetting occurs because the gradient update of the current task does not consider generalizability among all tasks. Restricting the updating of vital weights of previous tasks, as the regularization-based method(Kirkpatrick et al., 2017) does, is the direct way to our goal. (Lin et al., 2022) points out that the deeper the layers, the more the memory footprint occupied by weights increases, while the size of feature maps becomes smaller and smaller. Our idea is that since both can reflect output patterns, we can memorize previous feature-extracting patterns by regulating important feature maps to overcome forgetting with a little extra consumption.
@@ 5.2.2. How to Select Reference
The intuitive idea is to regulate changes of vital feature maps produced by each sample of previous tasks during training for a new task. However, it is impossible to store samples continuously in resource-limited scenarios. As the number of tasks increases, so does the number of samples required. Alternatively, we regulate feature maps from the current sample instead of previous samples. That’s because regulating output patterns of feature maps is unrelated to the information of samples. The only effort is to recognize feature maps that require regulation. To realize this, we identify locations of feature maps vital to previous tasks by $\mathscr{l}_{1}$-norm(Li et al., 2017) with ratio $R$.
@@ 5.2.3. Detailed Process
After training the (t-1)-th task, neural network $\theta^{t-1}$ is generated. We get important positions of feature maps from the first batch of (t-1)-th task. Assume $(i,j)$ denotes j-th feature map in i-th layer to be position, all important positions are recorded in a set $I$. Before training new t-th task, we first select a few samples, defined as $\mathcal{M}$, from the new task to run it on the neural network $\theta^{t-1}$ and get feature map standard, called $FS$, to be referenced. The process is defined as $FS=F(\theta^{t-1},\mathcal{M},I)$, where $F$ is the function to get feature maps of important position $I$. Note that we only get $FS$ before training the new task without keeping acquiring during training, shown in the gray background part at the bottom of the Figure  3.
During training the t-th task, we integrate $M$ in each batch and get the corresponding feature maps. $FM$ is calculating from $FM=F(\theta^{t},\mathcal{M},I)$. The limiting changes of vital feature maps between $FS$ and $FM$ can be described as the regulation loss function
@@
The entire loss function $\mathcal{L}$ is the combination of cross-entropy loss function $\mathcal{L}_{c}$ and the regulation loss $\mathcal{L}_{f}$, described as follow
@@
After the t-th task, we still use the first batch to get the important position $I^{\prime}$ of the current task with ratio $R$ and combine it with the last version, $I=I\lor I^{\prime}$. The whole process of the algorithm is shown in Algorithm 1. As shown in Section 6.3, Memorize Feature Patterns successfully delays forgetting with a few extra resources. The combination of Maintaining Generalizability and Memorize Feature Patterns outperforms other state-of-the-art   CL  methods(see in Table 1). Different from other replay methods, we are not required to store samples in storage, as they are only utilized during the current training. This also enhances privacy security.
Input :  Important Position Set $I$; Mem Sample $\mathcal{M}$; Function of Getting Feature Map $F$; Function of Neural Network $f$; Cross-entropy loss $\mathcal{L}_{c}$; Feature Regularization $\mathcal{L}_{f}$; Selecting ratio $R$
@@ 6. Experiment
Table 1. Comparison with CL  methods on pre-trained model. Classification results on ResNet-18 for standard CL  benchmarks, averaged across $5$ runs. LightCL  is outstanding beyond other CL  methods with fewer resources.
Method Buffer Sparsity Split CIFAR10 Split TinyImageNet
@@
‘/’ indicates the corresponding item is not needed (Buffer is not required in JOINT, SGD, EWC, PNN) or cannot run (PNN cannot run
@@
Table 2. Comparison with CL  methods from scratch. LightCL  still outperforms other CL  methods.
Method Split CIFAR10 Split TinyImageNet
@@ 6.1. Experiment Setting
Datasets: We evaluate LightCL  on Split CIFAR-10(Krizhevsky et al., 2009) and Split Tiny-ImageNet(Deng et al., 2009). To compare with other CL  methods, we strictly follow (Buzzega et al., 2020; Wang et al., 2022) by splitting CIFAR-10 and Tiny-ImageNet into $5$ and $10$ tasks, and each task including $2$ and $20$ classes, respectively. For Split CIFAR-10 and Tiny-ImageNet, we train each task for $50$ and $100$ epochs with a batch size of $32$.
@@
Baseline Mehtods: We select typical CL  methods from three categories, mentioned in section 2, including regularization-based (EWC(Kirkpatrick et al., 2017)), replay (DER(Buzzega et al., 2020), ER(Chaudhry et al., 2019), FDR(Benjamin et al., 2019)), and architecture-based (PNN(Rusu et al., 2016)) methods. Since these above methods do not consider the edge scenario, we also compare our method against the state-of-the-art   efficient CL  method SparCL(Wang et al., 2022) combined with the replay method DER++ and ER. We use SGD without any CL  method as a lower bound and train all tasks jointly (JOINT) as an upper bound.
@@
Metrics: We use AA, mentioned in 3, to evaluate the performance of overcoming forgetting. Moreover, we follow (Wang et al., 2022) to theoretically assess the training FLOPs and peak memory footprint (including feature maps, gradients, and model parameters during training) for the illustration of resource usage.
@@
Setting of LightCL: We use a pre-trained ResNet-18 model (pre-trained on the ImageNet32323232$\times$$32$(Chrabaszcz et al., 2017)), and freeze all lower and middle layers (before Layer4.0.conv1 in ResNet-18). For regulating the feature maps, we select vital feature maps with a ratio $R=15$% in each layer after Layer4.0.conv1 (included). Hyperparameter $\beta$ in Equation 4 equals $0.0002$. Although our method doesn’t need to store samples in storage, for a fair comparison, we set the same buffer size to compare LightCL  with replay methods.
@@
Experiment Details: We operate our experiments on a server equipped with an NVIDIA A800 GPU in Section 6.2 and Section 6.3. We further verify our method’s effectiveness on an edge device, Jetson Nano, in Section 6.4. All methods employ ResNet-18(He et al., 2016) with a learning rate of $0.01$. We set other hyperparameters as suggested in (Buzzega et al., 2020; Wang et al., 2022) for other CL  methods. For a fair comparison, all other methods employ the same pre-trained model in Section 6.2.1. Although the pre-trained model is more practical for application, most of CL  methods don’t take this into account and train the model from scratch. Hence, we test all methods without the pre-trained model to evaluate the effectiveness of LightCL  in Section 6.2.2. In order to compare with SparCL, we prune our model with sparsity ratio $90\%$ as the same as SparCL. Specifically, LightCL  uses static filter pruning for ResNet-18 and sets sparsity ratio as follows: 1) before Layer3.0 as $20\%$; 2) Layer3.0 and Layer3.1 as $70\%$; 3) others as $80\%$. Sparse is not the key point in our paper, and we set it up this way because we want to retain enough knowledge in the frozen layers.
@@ 6.2. Main Result
We evaluate results on Split CIFAR-10 and Tiny-ImageNet with a pre-trained ResNet-18 model under both CIL and TIL settings. The compared methods are divided into normal CL  methods (EWC, PNN, FDR, ER, DER++) and an efficient CL  method (SparCL). As shown in Table 1, the accuracy of LightCL  without sparsity increases compared with ER, which performs best in normal CL  methods except PNN. On Split Tiny-ImageNet, FLOPs of LightCL  greatly reduce at most $3.07$$\times$ and Memory footprint at most $4.79$$\times$, compared with normal CL  methods. That’s because LightCL  reduces the computation of backward process, and memory footprint of gradient and intermediate feature maps in frozen layers. Although the architecture-based method PNN has excellent accuracy under the TIL setting, it has tremendous FLOPs and memory footprint usage and also ignores the real application always under the CIL setting where the task’s identity is not known during testing. On the other hand, compared with the efficient method SparCL, the delaying performance of LightCL  with the same sparsity ratio is still competitive. Note that we employ a simple sparsity strategy without any optimization, such as dynamic sparsity(Wang et al., 2022). Surprisingly, the memory footprint is reduced at most $6.19$$\times$ compared with SparCLDER++. Although the FLOPs of LightCL  is higher than SparCL, LightCL  is faster than SparCL in the real application since SparCL still costs abundant time on memory transfer. This will be discussed in Section 6.4.
@@
Although deploying a pre-trained model on edge devices is more practical, most CL  methods ignore it and focus more on CL  from scratch. To better illustrate the effectiveness of LightCL, we also compare LightCL  with other CL  methods in the setting of CL  from scratch. Since LightCL  needs to freeze the lower and middle layers, we take the first task in the dataset as a pre-trained process, and we freeze the lower and middle layers in the following tasks. From Table 2, LightCL  is still outstanding compared with other methods. LightCL0.9 with sparsity ratio $90\%$ also shows better performance compared with the SparCL. Although replay methods (FDR, ER, DER++) are ineffective with a limited buffer size of $15$, LightCL  is favorable in a limited buffer size.
@@ 6.3. Ablation
In this section, we illustrate the effectiveness of Maintaining Generalizability and Memorize Feature Patterns of LightCL  in Figure  4.
Discussion of Maintaining Generalizability: Maintaining Generalizability is an easy but effective method to delay forgetting and reduce resource usage. From Figure  4, compared with red and blue bars, all CL  methods can enhance performance with the pre-trained model. Most of them (SGD, EWC, FDR, ER, LightCL) can even further delay the forgetting problem by Maintaining Generalizability. Note that the accuracy of SparCL decreases when using Maintaining Generalizability. This phenomenon is because frozen layers weaken the dynamic pruning process, leading to the accuracy drop. Although most of CL  methods can benefit from Maintaining Generalizability, LightCL  still performs better than other methods.
Figure 4. Ablation experiments on Split CIFAR10 dataset under the TIL setting across 5 runs. Red bars denote the setting without the pre-trained model and Maintaining Generalizability (freezing layers). Blue bars denote the setting only with the pre-trained model. Green bars have both of them.
Discussion of Memorize Feature Patterns: We compare the performance of LightCL  with and without Memorize Feature Patterns. As shown in Figure  4, the green bars of SGD and LightCL  are the LightCL  with and without Memorize Feature Patterns, respectively. We can see that LightCL  with Memorize Feature Patterns can further increase the accuracy, proving its effectiveness. Let’s consider the resource consumption of this method. We need to store the Feature Standard $FS$, mentioned in Section 5.2, as a reference to regulate. Since we only need to store important feature maps after Layer4.0.conv1 (included), the overall consumption of memory footprint, suppose all feature maps need to be stored, occupies about $1.9$MB (on Split CIFAR-10), which is pretty small compared to what we reduce by Maintaining Generalizability.

### v3

@@ 5. Proposed Architecture
LightCL, as shown in Figure  3, efficiently delays forgetting through Maintain Generalizability and Memorize Feature Patterns. All these methods are based on the discovery of generalizability in Section 4. We first introduce Maintain Generalizability in Section 5.1 and then illustrate Memorize Feature Patterns in Section 5.2.
Figure 3. Illustration of each component in LightCL. Before training a new task, we obtain feature map standards as references for later regulation and store them in memory, as shown in the bottom part of the figure. During the training process, shown in the top part of the figure, we Maintain Generalizability by freezing lower and middle layers, and Memorize Feature Patterns by regulation between current feature maps and stored feature map standards.
5.1. Maintain Generalizability
@@
According to the discovery in Section 4 that lower and middle layers have strong generalizability, the training consumption on them can be a kind of redundancy. Hence, we easily freeze them to maintain generalizability without the training process, illustrated in the gray background section at the top of the Figure  3. Specifically, we must pretrain a network, freeze its lower and middle layers, and finally deploy it on the edge device for CL. Training the t-th task can be formulated as
@@
(3) $\begin{array}[]{cl}\theta_{i}^{t}\leftarrow\theta_{i}^{t}+\alpha\nabla_{\theta_{i}^{t}}\mathcal{L}_{c}(f(\theta^{t},x_{t}),y_{t})\quad\text{s.t.}\quad i>\textrm{FZ},\end{array}$
@@
where FZ is the last layer index of frozen layers, $\mathcal{L}_{c}$ is the cross-entropy loss function, and function $f$ represents forward pass. For example, we set FZ as layer3.1.conv2 based on the discovery when we utilize ResNet-18 as the backbone network. This setting will be kept for experiments in Section 6.
Maintain Generalizability greatly reduces resource consumption. Firstly, as frozen layers do not participate in backpropagation, there is no need to compute and store gradients and feature maps during training, which greatly reduces memory usage. Secondly, the complexity of the computation graph shrinks, so backpropagation requires less computation. This alleviates the burden on hardware, thereby improving computational efficiency and shortening training time. Lastly, frozen layers decrease data transfer in memory during backward propagation, hence reducing latency and improving data flow efficiency. Notably, this high-efficiency method is able to delay the forgetting procedure without diminishing the knowledge of counterparts. As evidenced by the experiments in Section 6.2.1, with the help of Maintain Generalizability, LightCL  reduces $6.16$$\times$ memory footprint and is competitive beyond the efficient CL  method SparCL in FLOPs. Meanwhile, it has greatly delayed forgetting by the seamless integration with most CL  methods in Section 6.3, which illustrates the powerful scalability of our method.
@@ 5.2.1. How to Memorize Efficiently
Forgetting occurs because the gradient update of the current task does not consider generalizability among all tasks. Restricting the updating of vital weights of previous tasks, as the regularization-based method (Kirkpatrick et al., 2017) does, is the direct way to achieve our goal. (Lin et al., 2022) points out that the deeper the layers, the more the memory footprint occupied by weights, while the size of feature maps becomes smaller and smaller. Our idea is that since both feature maps and weights can reflect the output patterns of the model, we can memorize previous feature-extracting patterns by regulating important feature maps to overcome forgetting with a little extra consumption.
@@ 5.2.2. How to Select Reference
The intuitive idea is to regulate changes in vital feature maps produced by each sample of previous tasks during training for a new task. However, it is impossible to store samples continuously in resource-limited scenarios. As the number of tasks increases, so does the number of samples required. Alternatively, we regulate feature maps from the current sample instead of previous samples because regulating output patterns of feature maps is unrelated to the information of samples. The only effort is to recognize feature maps that require regulation. To realize this, we identify positions of feature maps vital to previous tasks by $\mathscr{l}_{1}$-norm (Li et al., 2017) with a ratio $R$.
@@ 5.2.3. Detailed Process
After training the (t-1)-th task, neural network $\theta^{t-1}$ is generated. We get important positions of feature maps from the first batch of (t-1)-th task. Assume position $(i,j)$ denotes j-th feature map in i-th layer. All important positions are recorded in a set $I$. Before training new t-th task, we first select a few samples, defined as $\mathcal{M}$, from the new task to run it on the neural network $\theta^{t-1}$ and get feature map standards, called $FS$, as references. The process is defined as $FS=F(\theta^{t-1},\mathcal{M},I)$, where $F$ is the function to get feature maps of important position $I$. Note that we only get $FS$ before training the new task without continually acquiring it during training, shown in the bottom of Figure  3.
During training the t-th task, we integrate $M$ in each batch and get the corresponding feature maps. $FM$ is calculated from $FM=F(\theta^{t},\mathcal{M},I)$. The limiting changes of vital feature maps between $FS$ and $FM$ can be described as the regulation loss function
@@
The entire loss function $\mathcal{L}$ is the combination of cross-entropy loss function $\mathcal{L}_{c}$ and the regulation loss $\mathcal{L}_{f}$, described as
@@
After the t-th task, we use the first batch to get the important position $I^{\prime}$ of the current task with a ratio $R$ and combine it with the last version, $I=I\lor I^{\prime}$. The whole process of the algorithm is shown in Algorithm 1. As shown in Section 6.3, Memorize Feature Patterns successfully delays forgetting with a few extra resources. The combination of Maintain Generalizability and Memorize Feature Patterns outperforms other state-of-the-art   CL  methods, shown in the Table 1. Different from other replay methods, we do not store samples in storage as they are only utilized during the current training. This also enhances privacy security.
Input : Important Position Set $I$; Mem Sample $\mathcal{M}$; Feature Standard $FS$; Feature Map $FM$; Function of Getting Feature Map $F$; Function of Neural Network $f$; Cross-entropy loss $\mathcal{L}_{c}$; Feature Regularization $\mathcal{L}_{f}$; Selecting ratio $R$
@@ 6. Experiment
Table 1. Comparison with other CL  methods on pre-trained model. Classification results on ResNet-18 for standard CL  benchmarks, averaged across $5$ runs. LightCL  is outstanding beyond other CL  methods with fewer resources.
Method Buffer Sparsity Split CIFAR-10 Split TinyImageNet
@@
‘/’ indicates the corresponding item is not needed (Buffer is not required in JOINT, SGD, EWC, and PNN) or cannot run (PNN cannot run
@@
Table 2. Comparison with CL  methods from scratch with ResNet-18 and ResNet-50. LightCL  still outperforms other CL  methods and shows outstanding performance on both large and small networks from scratch.
Method Split CIFAR-10 Split TinyImageNet
@@ 6.1. Experiment Setting
Datasets. We evaluate LightCL  on Split CIFAR-10 (Krizhevsky et al., 2009) and Split Tiny-ImageNet (Deng et al., 2009). To compare with other CL  methods, we strictly follow (Buzzega et al., 2020; Wang et al., 2022) by splitting CIFAR-10 and Tiny-ImageNet into $5$ and $10$ tasks, and each task including $2$ and $20$ classes, respectively. For Split CIFAR-10 and Tiny-ImageNet, we train each task for $50$ and $100$ epochs with a batch size of $32$.
@@
Baseline Mehtods. We select typical CL  methods from three categories mentioned in section 2, including regularization-based (EWC (Kirkpatrick et al., 2017)), replay (DER (Buzzega et al., 2020), ER (Chaudhry et al., 2019), and FDR (Benjamin et al., 2019)), and architecture-based (PNN (Rusu et al., 2016)) methods. Since these above methods do not consider the edge scenario, we also compare our method against the state-of-the-art   efficient CL  method SparCL (Wang et al., 2022) combined with the replay method DER++ and ER. We use SGD without any CL  method as a lower bound and train all tasks jointly (JOINT) as an upper bound.
@@
Metrics. We use AA, mentioned in Section 3, to evaluate the performance of overcoming forgetting. Moreover, we follow (Wang et al., 2022) to theoretically assess the training FLOPs and peak memory footprint (including feature maps, gradients, and model parameters during training) for the illustration of resource usage.
@@
Setting of LightCL. We use a pre-trained ResNet-18 model (pre-trained on the ImageNet32323232$\times$$32$ (Chrabaszcz et al., 2017)), and freeze all lower and middle layers (before Layer4.0.conv1 in ResNet-18). For regulating the feature maps, we select vital feature maps with a ratio $R=15$% in each layer after Layer4.0.conv1 (included). Hyperparameter $\beta$ in Equation 4 equals $0.0002$. Although our method does not need to store samples in storage, we set the same buffer size to compare LightCL  with replay methods for a fair comparison.
@@
Experiment Details. We operate our experiments on a server equipped with an NVIDIA A800 GPU in Section 6.2 and Section 6.3. We further verify the practical effectiveness of our method on an edge device, Jetson Nano, in Section 6.4. All methods employ ResNet-18 (He et al., 2016) with a learning rate of $0.01$. We set other hyperparameters as suggested in (Buzzega et al., 2020; Wang et al., 2022) for other CL  methods. For a fair comparison, all other methods employ the same pre-trained model in Section 6.2.1. Although the pre-trained model is more practical for application, most of CL  methods do not take this into account and pay more attention to training the model from scratch. Hence, we test all methods without the pre-trained model to evaluate the effectiveness of LightCL  in Section 6.2.2. In order to compare with SparCL, we prune our model with sparsity ratio $90\%$ as the same as SparCL. Specifically, LightCL  uses static filter pruning for ResNet-18 and sets sparsity ratio as follows: 1) before Layer3.0 as $20\%$; 2) Layer3.0 and Layer3.1 as $70\%$; 3) others as $80\%$. Sparse is not the key point in our paper, and we set it up this way because we want to retain enough knowledge in the frozen layers.
@@ 6.2. Main Result
We evaluate results on Split CIFAR-10 and Tiny-ImageNet with a pre-trained ResNet-18 model under both CIL and TIL settings. The compared methods are divided into normal CL  methods (EWC, PNN, FDR, ER, and DER++) and an efficient CL  method (SparCL). As shown in Table 1, the accuracy of LightCL  without sparsity is higher than ER, which performs best in normal CL  methods except PNN. On Split Tiny-ImageNet, FLOPs of LightCL  greatly reduce at most $3.07$$\times$ and Memory footprint at most $4.79$$\times$, compared with normal CL  methods. That is because LightCL  reduces the computation of backward process, and memory footprint of gradient and intermediate feature maps in frozen layers. However, other algorithms significantly increase resource consumption to mitigate forgetting. For example, EWC requires storing twice the weight size because it calculates the importance of each weight. In the Replay method (FDR, ER, and DER++), during training a new task, previous samples from the buffer need to be included in each batch, which greatly increases memory and computation consumption. Among them, DER++ is particularly severe as it needs to calculate twice the buffer size, leading to even greater increases. Although the architecture-based method PNN has excellent accuracy under the TIL setting, it has tremendous FLOPs and memory footprint usage since it requires an additional structure for a new task and its forward process requires all structures during training. It also ignores the real application always under the CIL setting where the identity of a certain task is not known during testing. On the other hand, compared with the efficient method SparCL, the delaying performance of LightCL  with the same sparsity ratio is still competitive. Note that we employ a simple sparsity strategy without any optimization, such as dynamic sparsity (Wang et al., 2022). Surprisingly, the memory footprint is reduced at most $6.19$$\times$ compared with SparCLDER++. Although the FLOPs of LightCL  is higher than SparCL, LightCL  is faster than SparCL in the real application since SparCL still costs abundant time on memory transfer. This will be discussed in Section 6.4.
@@
Although deploying a pre-trained model on edge devices is more practical, most CL  methods ignore it and focus more on CL  from scratch. To better illustrate the advantages of LightCL, we also compare LightCL  with other CL  methods in the setting of CL  from scratch. To further demonstrate the performance of LightCL  on a large network, we compare LightCL  with outstanding CL methods (ER and DER++) on ResNet-50. Since LightCL  needs to freeze the lower and middle layers, we take the first task in the dataset as a pre-trained process, and we freeze the lower and middle layers in the following tasks. For simplicity, the experimental settings of ResNet-50 are the same as ResNet-18 and we set frozen layers before layer4.0.conv1 in ResNet-50. From Table 2, LightCL  is still outstanding compared with other methods in both ResNet-18 and ResNet-50. When using ResNet-18, LightCL0.9 with sparsity ratio $90\%$ also shows better performance compared with the SparCL. Although replay methods (FDR, ER, and DER++) are ineffective with a limited buffer size of $15$, LightCL  is favorable in a limited buffer size.
@@ 6.3. Ablation
In this section, we illustrate the effectiveness of Maintain Generalizability and Memorize Feature Patterns of LightCL  with ablation experiments in Figure  4.
Discussion of Maintain Generalizability. Maintain Generalizability is an easy but effective method to delay forgetting and reduce resource usage. From Figure  4, compared with red and blue bars, all CL  methods can enhance performance with the pre-trained model. Most of them (SGD, EWC, FDR, ER, and LightCL) can even further delay the forgetting problem by Maintain Generalizability, shown in green bars. Note that the accuracy of SparCL decreases when using Maintain Generalizability. This phenomenon occurs because frozen layers weaken the dynamic pruning process, leading to a drop in accuracy. Although most of CL  methods can benefit from Maintain Generalizability, LightCL  still performs better than other methods.
Figure 4. Ablation experiments on Split CIFAR-10 dataset under the TIL setting across 5 runs. Red bars denote the setting without the pre-trained model and Maintain Generalizability (freezing layers). Blue bars denote the setting only with the pre-trained model. Green bars have both of them.
Discussion of Memorize Feature Patterns. We compare the performance of LightCL  with and without Memorize Feature Patterns. As shown in Figure  4, the green bars of SGD and LightCL  are able to represent the LightCL  with and without Memorize Feature Patterns, respectively. The result shows that LightCL  with Memorize Feature Patterns can further increase the accuracy, proving its effectiveness. Considering the resource consumption of this method, we need to store the Feature Standard $FS$, mentioned in Section 5.2, as a reference to regulate. As we only need to store important feature maps after Layer4.0.conv1 (included), the overall consumption of memory footprint, suppose all feature maps need to be stored, occupies about $1.9$MB (on Split CIFAR-10), which is pretty small compared to what we reduce by Maintain Generalizability.